# arXiv:2603.03190v1[cs.AI] 03 Mar 2026

### v1

@@ Methods
where $y$ denotes the ground-truth Song ID and $\mathrm{CE}(\cdot)$ represents the cross-entropy applied to the logits. This branch is introduced as an auxiliary objective during pretraining, following the design philosophy of SupMAE [supmae].
@@
This formulation follows the core idea of SupMAE: combining a masked generative objective with a supervised classification branch [supmae]. In PredANN++, masked teacher prediction enforces local, time-resolved alignment between EEG and stimulus-derived representations, while the classification objective promotes global semantic organization of EEG representations.
Relation to prior works. Compared with PredANN, which is built upon a contrastive CNN architecture optimized via an InfoNCE objective, PredANN++ adopts a Transformer-based generative masked modeling paradigm. This change represents a shift from contrastive alignment toward structured generative pretraining, following the design principles that have proven effective in masked autoencoding frameworks across multiple modalities [devlin-etal-2019-bert, mae, beit, mae_audio1, jiang2024large].
@@
Multitask pretraining was conducted for 10,000 epochs under a unified experimental setting. All multitask models were trained with a fixed random seed of 42 to ensure reproducibility. Optimization was performed using Adam for both pretraining and fine-tuning stages, with a learning rate of 0.003 and a batch size of 48. The checkpoint obtained at the final pretraining epoch was used as the initialization for subsequent fine-tuning. Across all teacher types (Acoustics, Surprisal, and Entropy), the EEG encoder architecture was kept identical, ensuring that differences in downstream performance arise solely from the choice of teacher representation rather than architectural variation.
@@
As a baseline, we trained the same encoder-only architecture for Song ID classification from scratch without multitask pretraining (Fullscratch). For a fair comparison, the Fullscratch models were trained for the same number of epochs as the fine-tuning stage, namely 3,500 epochs, using identical optimization settings. To quantify the effect of initialization diversity, Fullscratch models were trained with multiple random seeds (e.g., 42, 0, 1, and 2). These independently trained models were further used to evaluate seed-based ensemble effects, providing a direct comparison with ensembles constructed from models pretrained with different teacher representations.
@@
At inference time, we evaluate the complementarity of learned EEG representations using an equal-weight Deep Ensemble strategy  [DeepEnsemble2017]. Let $K$ denote the number of independently trained classifiers included in the ensemble, with $K\in\{2,3\}$. Given an EEG input $\mathbf{x}$, each classifier $f^{(k)}$ outputs a vector of logits
@@
The Naturalistic Music EEG Dataset – Tempo (NMED-T) [losorelli2017nmed] analyzed in this study is publicly available at: https://exhibits.stanford.edu/data/catalog/jn859kj8079.
@@
The complete source code for dataset preprocessing, model training, evaluation, and implementation of the proposed method is publicly available at https://github.com/ShogoNoguchi/PredANNpp. The repository provides all scripts necessary to reproduce the experiments, including multitask pretraining, finetuning, evaluation pipelines, and a Gradio-based inference demo. The code is released under the CC-BY-SA 4.0 license.

### v2

@@ Methods
where $y$ denotes the ground-truth Song ID and $\mathrm{CE}(\cdot)$ represents the cross-entropy applied to the logits. This branch is introduced as an auxiliary objective during pretraining, following the design philosophy of SupMAE [55].
@@
This formulation follows the core idea of SupMAE: combining a masked generative objective with a supervised classification branch [55]. In PredANN++, masked teacher prediction enforces local, time-resolved alignment between EEG and stimulus-derived representations, while the classification objective promotes global semantic organization of EEG representations.
Relation to prior works. Compared with PredANN, which is built upon a contrastive CNN architecture optimized via an InfoNCE objective, PredANN++ adopts a Transformer-based generative masked modeling paradigm. This change represents a shift from contrastive alignment toward structured generative pretraining, following the design principles that have proven effective in masked autoencoding frameworks across multiple modalities [65, 68, 69, 73, 71].
@@
Multitask pretraining was conducted for 10,000 epochs under a unified experimental setting. All multitask models were trained with a fixed random seed of 42 to ensure reproducibility. Optimization was performed using Adam [74] for both pretraining and fine-tuning stages, with a learning rate of 0.003 and a batch size of 48. The checkpoint obtained at the final pretraining epoch was used as the initialization for subsequent fine-tuning. Across all teacher types (Acoustics, Surprisal, and Entropy), the EEG encoder architecture was kept identical, ensuring that differences in downstream performance arise solely from the choice of teacher representation rather than architectural variation.
@@
As a baseline, we trained the same encoder-only architecture for Song ID classification from scratch without multitask pretraining (Fullscratch). For a fair comparison, the Fullscratch models were trained for the same number of epochs as the fine-tuning stage, namely 3,500 epochs, using identical optimization settings. To quantify the effect of initialization diversity, Fullscratch models were trained with multiple random seeds (e.g., 0, 1, and 42). These independently trained models were further used to evaluate seed-based ensemble effects, providing a direct comparison with ensembles constructed from models pretrained with different teacher representations.
@@
At inference time, we evaluate the complementarity of learned EEG representations using an equal-weight Deep Ensemble strategy  [44]. Let $K$ denote the number of independently trained classifiers included in the ensemble, with $K\in\{2,3\}$. Given an EEG input $\mathbf{x}$, each classifier $f^{(k)}$ outputs a vector of logits
@@
The Naturalistic Music EEG Dataset – Tempo (NMED-T) [41] analyzed in this study is publicly available at: https://exhibits.stanford.edu/data/catalog/jn859kj8079.
@@
The complete source code for dataset preprocessing, model training, evaluation, and implementation of the proposed method is publicly available at https://github.com/ShogoNoguchi/PredANNpp. The repository provides all scripts necessary to reproduce the experiments, including multitask pretraining, finetuning, evaluation pipelines, and a Gradio-based inference demo. Furthermore, we provide a comprehensive project page at https://shogonoguchi.github.io/PredANNpp/, which serves as an online supplement providing the project overview, results tables, pretrained checkpoints, interactive audio-synchronized feature visualizations, Gradio demo videos, and dataset information. The code is released under the CC-BY-SA 4.0 license.
Supplementary information
Supplementary Note 1: Algorithmic details of MuQ extraction
In Algorithm~1, the input waveform is transformed explicitly as \$x x_1 x_2 x_3\$. In Algorithm~2, K-means is run with K-means++ initialization, random_state=0, and n_init=10.
Algorithm 1 Extraction of frame-wise continuous MuQ embeddings from 30-second audio
1:Audio waveform \$x\$ and pre-trained MuQ model with frozen parameters \$ϕ\$
2:Frame-wise continuous MuQ representation \$M_raw = (m_raw_t)_t=1^750\$, where \$m_raw_t R^1024\$
3:Convert \$x\$ to mono and resample to \$24~kHz\$ to get \$x_1\$
4:Trim or zero-pad \$x_1\$ to exactly \$30\$ seconds to get \$x_2\$ $\triangleright$ MuQ expects fixed-length 30-s inputs
5:Normalize \$x_2\$ according to MuQ preprocessing to get \$x_3\$
6:Obtain MuQ hidden states \$H\$ from \$MuQ_ϕ(x_3)\$
7:Extract the final-layer representation \$H_last R^1 ×750 ×1024\$ from \$H\$
8:for \$t 1\$ to \$750\$ do
9:  \$m_raw_t H_last,1,t,:\$
10:end for
11:return \$M_raw\$
Algorithm 2 K-means discretization of frame-wise MuQ embeddings
1:Pooled MuQ frame embeddings \$E = _songs _chunks {m_raw_t}\$ and number of clusters \$K = 128\$
2:For each 30-second chunk \$c {1,…,C}\$, a discrete MuQ sequence \$M^(c)_disc = (m^(c)_disc_t)_t=1^750\$, where \$m^(c)_disc_t {0,…,K-1}\$
3:Fit K-means to \$E\$ with \$K\$ clusters
4:Obtain cluster centroids \${c_k}_k=0^K-1\$
5:for each 30-second chunk \$c\$ do
6:  for \$t 1\$ to \$750\$ do
7:    \$m^(c)_disc_t ←arg min_k {0,…,K-1} m^(c)_raw_t - c_k _2\$
8:  end for
9:end for
10:return \${M^(c)_disc}_c=1^C\$
Supplementary Note 2: Algorithmic details of computing Surprisal, Entropy, and discretization
For Algorithm~3, we write \$s^(j)_raw = (s^(j)_raw,t)_t=1^150\$ and \$h^(j)_raw = (h^(j)_raw,t)_t=1^150\$ for the 3-second Surprisal and Entropy sequences of segment \$j\$. For Algorithm~4, we write \$s^(j)_disc = (s^(j)_disc,t)_t=1^150\$ and \$h^(j)_disc = (h^(j)_disc,t)_t=1^150\$ for the corresponding discretized sequences. Here, \$BinIndex(u;E)\$ returns the unique bin index \$b {0,…,B-1}\$ such that \$E_b u < E_b+1\$, with the maximum value assigned to bin \$B-1\$.
Algorithm 3 Sliding-window computation of frame-wise Surprisal and Entropy using MusicGen k1
1:Full-length audio waveform \$x\$ and context window length \$W {8,16,32}\$ seconds
2:Segment-wise continuous predictive representations \$S_raw = {s^(j)_raw}_j=0^N_seg-1\$ and \$H_raw = {h^(j)_raw}_j=0^N_seg-1\$
3:Convert \$x\$ to mono and resample to \$32~kHz\$ to get \$x_1\$
4:Encode \$x_1\$ using EnCodec to obtain \$z N^4 ×T_frames\$ $\triangleright$ 4 codebooks at 50~Hz
5:\$L 150\$ $\triangleright$ 3-s segment length in frames
6:\$d 5\$ $\triangleright$ 0.1-s stride in frames
7:\$N_seg ←⌊T_frames - Ld + 1\$
8:for \$j 0\$ to \$N_seg-1\$ do
9:  \$s_j j d\$
10:  \$e_j s_j + L\$
11:  \$W_f 50 W\$
12:  Construct the context token sequence from frames \$[e_j - W_f,  e_j)\$ to get \$z^(j)_ctx\$
13:  if \$e_j - W_f < 0\$ then
14:    Left-pad the missing prefix of \$z^(j)_ctx\$ with the MusicGen special token
15:  end if
16:  Compute k1 logits from MusicGen conditioned on \$z^(j)_ctx\$ under the empty-text condition
17:  Extract the k1 logits corresponding to the final \$L\$ frames
18:  for \$t 1\$ to \$L\$ do
19:    \$s^(j)_raw,t -logp_θ(z_s_j+t C_j,t)\$
20:    \$h^(j)_raw,t -_v p_θ(v C_j,t)logp_θ(v C_j,t)\$
21:  end for
22:end for
23:return \$S_raw\$, \$H_raw\$
Algorithm 4 Quantile-based discretization of Surprisal and Entropy
1:Segment-wise continuous representations \$S_raw\$ and \$H_raw\$, and number of bins \$B = 128\$
2:Segment-wise discrete representations \$S_disc = {s^(j)_disc}_j=0^N_seg-1\$ and \$H_disc = {h^(j)_disc}_j=0^N_seg-1\$
3:Pool all Surprisal values across all segments into \$S\$
4:Pool all Entropy values across all segments into \$H\$
5:for \$k 0\$ to \$B\$ do
6:  \$q_k k / B\$
7:  \$E^(S)_k ←Quantile(S, q_k)\$
8:  \$E^(H)_k ←Quantile(H, q_k)\$
9:end for
10:for each segment \$j\$ do
11:  for \$t 1\$ to \$150\$ do
12:    \$s^(j)_disc,t ←BinIndex(s^(j)_raw,t; E^(S))\$
13:    \$h^(j)_disc,t ←BinIndex(h^(j)_raw,t; E^(H))\$
14:  end for
15:end for
16:return \$S_disc\$, \$H_disc\$
Supplementary Note 3: Qualitative inspection of context-dependent Surprisal/Entropy time courses
Supplementary Figure~12 provides an example visualization in which MusicGen-derived Surprisal and Entropy are displayed together with acoustic representations and human EEG from Subject~2 on a common time axis.
In this visualization, scalp EEG signals recorded with the EGI 128-channel HydroCel Geodesic Sensor Net are displayed using stacked traces for all available channels. This representation allows the temporal dynamics of large-scale scalp EEG activity to be inspected together with the acoustic and model-derived features on the same time axis.
Two qualitative observations are worth noting. First, a consistent qualitative trend across context lengths can be observed: when the context window increases from 8~s to 16~s and 32~s, the Surprisal and Entropy trajectories appear progressively smoother and vary on slower time scales. This visual smoothing is compatible with the idea that longer contexts integrate information over longer musical histories, and therefore the resulting predictive quantities may show reduced sensitivity to short-term fluctuations.
Second, changes in the MusicGen-derived Surprisal and Entropy time courses often appear to co-occur with acoustically salient events visible in the acoustic panels, such as peaks in the RMS envelope and onset-dense regions in the mel-spectrogram. In particular, near RMS peaks and near-vertical broadband structures in the mel-spectrogram (which may correspond to percussive onset components), Surprisal and Entropy tend to show larger fluctuations. This behaviour is intuitively plausible: predictive quantities such as surprisal or entropy would naturally be expected to fluctuate more around strong acoustic events or percussive onsets. Therefore, the observed alignment provides qualitative face validity for our Surprisal/Entropy computation. Consistent with this interpretation, listening to Song~21 suggests that prominent RMS increases correspond to downbeat, low-frequency kick-like percussive pulses (sometimes accompanied by metallic pluck-/click-like transients).
For clarity, the mel-spectrogram panel titled ``Audio Mel Spectrogram (dB) | n_mels=128 mel_scale=htk'' is computed by applying a 128-bin mel filterbank (HTK mel scale) to the power spectrogram and converting the resulting mel-band power to decibels using \$10log_10()\$ for visualization. The RMS panel is computed as a short-time root-mean-square envelope of the waveform and is min--max normalized to the range \$[0,1]\$ within the plotted 30-second window.
To facilitate qualitative inspection of how predictive quantities relate to the acoustic structure of the music, we provide an interactive web-based visualization at https://shogonoguchi.github.io/PredANNpp/#syncviz. This interface enables synchronized playback of the audio signal together with its RMS envelope and the corresponding MusicGen-derived Surprisal and Entropy trajectories. The audio examples used in this visualization are drawn from the MTG-Jamendo dataset [47].
Figure 10: Performance of the conservative chunk-based computation.
Panels A--D compare Surprisal and Entropy computed independently within 30-s audio chunks (see Methods, ``Conservative chunk-based computation'') with those obtained using standard context-length settings and seed baselines.
A Classification accuracy using Surprisal features. Grey bars denote the three standard computations (context lengths: 8~s, 16~s, 32~s); the orange bar indicates the conservative 30-s chunk-based computation. Chunk-based Surprisal yields accuracy comparable to the best-performing 16~s context. B Classification accuracy using Entropy features. Grey bars denote the three standard computations (context lengths: 8~s, 16~s, 32~s); the green bar indicates the conservative 30-s chunk-based computation. In contrast to Surprisal (panel A), chunk-based Entropy does not reach the performance of the 8~s or 16~s contexts, although it remains above the 32~s condition. C Comparison of three no-pretraining seed baselines (Seed 0/1/42, grey) with conservative Surprisal (orange) and Entropy (green). Surprisal exceeds all seed baselines, whereas Entropy surpasses Seed 1 but does not exceed Seed 0 or Seed 42. D McNemar's test heatmap comparing each seed column with each conservative feature row. Blue cells indicate feature superiority, light grey indicates no significant difference; asterisks indicate significance levels (** \$p < 0.01\$, *** \$p < 0.001\$). Conservative Surprisal shows significant improvements over all seed baselines, whereas conservative Entropy reaches significance for Seed 1 only, with no significant differences observed for Seed 0 or Seed 42.
Figure 11: Validation accuracy across training epochs for MuQ, Surprisal, and Entropy representations when used directly for Song ID classification.
Each panel shows validation accuracy (%) as a function of training epoch for models that directly use discretized music-feature sequences as input, without EEG signals. For each 3-second segment, the input consists solely of discretized feature tokens: MuQ (75 steps at 25~Hz), Surprisal (150 steps at 50~Hz), or Entropy (150 steps at 50~Hz). Each scalar discrete value (0--127) is projected to a 512-dimensional embedding through a linear layer (1\$\$512). A learnable [CLS] token is prepended, and learnable positional embeddings are added. The sequence is processed by a 2-layer Transformer encoder (embedding dimension 512, 8 attention heads, MLP ratio 4.0, GELU activation, dropout 0.1), identical to the Transformer architecture used in the main EEG recognition experiments. The hidden representation of the [CLS] token is passed through a classification head composed of a linear layer (512\$\$256), batch normalization, ReLU activation, and a final linear layer (256\$\$10) to predict Song ID. The batch size is fixed to \$B=48\$. Optimization is performed using Adam. The learning rate is 0.003 for MuQ-based modeling and 0.0003 for Surprisal- and Entropy-based modeling, consistent with the learning-rate settings used in the main experiments. All curves represent exponential moving averages of validation accuracy. Under this modeling setup, all three representations independently achieve validation accuracies approaching 100%, demonstrating that each representation alone contains sufficient information to almost perfectly solve the 10-class Song ID task.
Figure 12: Example multi-panel visualization for Song 21 (30--60~s), Subject 2.
Panels show (top to bottom) scalp EEG from Subject~2 recorded with the EGI 128-channel HydroCel Geodesic Sensor Net (all channels stacked), the normalized short-time RMS envelope, the mel-spectrogram in decibels (128 mel bins; HTK mel scale), and MusicGen-k1-derived Surprisal and Entropy trajectories discretized into Q128 bins for context windows of 8, 16, and 32 seconds (see Supplementary Note 3: Qualitative inspection of context-dependent Surprisal/Entropy time courses for qualitative inspection).
Author information
Authors and Affiliations
Sony Computer Science Laboratories, Tokyo, Japan
Shogo Noguchi, Taketo Akama, Tai Nakamura, Shun Minamikawa & Natalia Polouliakh
Contributions
S.N. and T.A. jointly conceptualized the study. T.A. proposed the initial framework and research direction. S.N. refined and operationalized the methodology, designed the model architecture and experimental protocol, and led the software implementation. S.N. and T.N. conducted the experiments with feedback from T.A. ; T.N. additionally contributed to visualization and statistical testing, and assisted with supporting code for experiments and analyses. S.M. helped organize and finalize the codebase. S.N. wrote the first draft of the manuscript with contributions from T.N. and T.A.; S.N. and T.N. created the figures and tables, with feedback from T.A. T.A. and N.P. reviewed and edited the manuscript. N.P. advised the research and organized the research project.
Corresponding author
Correspondence to Shogo Noguchi and Taketo Akama.
Competing interests
The authors declare no competing interests.